OpenStudy (anonymous):

write an equation for: pays one cent for the first hour and doubles every subsequent hour.(for instance, you earn two cents the second hour, four cent for the 3rd hour , eight cent for the 4th and so forth) HELP PLEASE.

OpenStudy (mathmate):

Amount per hour, x=number of hours since start y=2^(x-1) Total amount y=2^x - 1

OpenStudy (anonymous):

thnx!

OpenStudy (mathmate):

You're welcome! :)